OpenStudy (anonymous):

Use the form y = mx + b for linear equation. {(x,y): (1,2), (2,5), (3,8),...} The answer is y = 3x - 1 I know how to get 3x it's by (5-2)/(2-1)= 3÷1=3 (8-5)/(3-2)= 3÷1=3 (8-2)/(3-1)= 6÷2=3 What I don't understand is how to get the -1

OpenStudy (freckles):

so you have y=3x+b to find b you need to take a point on y=mx+b say (1,2) y=3x+b replace x with 1 and replace y with 2 2=3(1)+b <--solve for b

OpenStudy (anonymous):

So then it would be 2 = 3(1) + b 3(1) = 3 so I'd have to carry 3 to the other side so its, -3 + 2 = b and I'd get -1 That solves b so you then write y = 3x -1? or did I do that wrong...

myininaya (myininaya):

that's right and (2,5) should work too y=3x+b 5=3(2)+b 5-6=b -1=b

myininaya (myininaya):

you see see for any point in that set you will get the b value to be -1

OpenStudy (anonymous):

Thank you! I actually understand it, not even that hard.